OpenStudy (anonymous):

Is 4x-3=19? true false open

OpenStudy (jwheele1):

one moment...

OpenStudy (anonymous):

false...... 4x-3 = 19 , add 3 to 19 and it is 22, 4x does not divide into 22 well

OpenStudy (anonymous):

x = 11/2 is true. x ≠ 11/2 is false.

OpenStudy (jwheele1):

Add three to both sides to isolate the variable 4x -3 = 19 +3 + 3 ----------------- 4x -0 = 22 OR 4x = 22 Now, divide both sides by the number glued to the variable, which is 4 4x = 22 --- --- 4 4 ___________ 1x = 5.5 OR x = 5.5 NOW that you know what "X" is, you can plug is back in to get a "balance statement" 4x -3 = 19 is now 4(5.5) -3 = 19 solve like normal (4 * 5.5) = 22 ... -3 = 19 SO the new problem looks like this (22 - 3) = 19 Then continue (22-3) = what? 19 ....=19 SO THEN is this statement true? 19 = 19? Does 19 equal 19? Yes. True.